APACHE POINT OBSERVATORY SDSS 2.5M OBSERVING LOG Friday November 24, 2006 (MJD 54064) ---=== OBSERVING TEAM ===--- Afternoon: Shannon Watters Night: Victor Malanushenko, Elena Malanushenko ---=== OBSERVING PLAN ===--- Science! ---=== OBSERVING SUMMARY ===--- We had heavy clouds during first part of night. Opened only at 6:45. We obtained Supernova 82-S stripe with lambda -156 to 122.5 including ramp with good photometric conditions. We also finished one plate 2386. We could not do spooling (see log). ---=== OBSERVING LOG ===--- Afternoon: --------- Cloudy skies. Temp ~7C, wind 15mph from the SSE. All systems checked and working. The spectrograph focus is set for ~5C. The MT Watcher is not running. Night: ----- 00:05Z Mounted imager inside enclosure 00:18Z Telescope bias run 6602 After bias ended we opened the doors. The dom is closed because of clouds and strong radar return 01:05 Close the doors and cloud cam because of rain sensor alarm, it was little snow even the humidity is 37% only. Rick called and requested 82-S stripe. 04:32 Door opened Watcher complained about id3 does not respond. Searching: im> iccSoundOff ........... ICC id3 is not available ( Bad ICC) ................... 04:50Z reboot of photo-1 crate. It helps to make id3 availble. 05:05Z Bias, run 6603 bad quality (stop button pressed inside enclosure) We did it to be sure that all systems works good after rebooting. We still have strong clouds and radar return and cannot open. Many times during the night watcher complaines about server connection, but fixes the problem itsef in several minuts. Several times I saw an error "im-serial" Camera Serial Line - server is not responding (Last packet: 2006-11-25 04:44:17Z), and many messages in murmurlog about broken sftp imager camera connection. I checked the server and did not find iop[im] telnet window. I rebooted the server and got all three expected iop windows, but did not get mt. We see radar return is moving away and we expect to clear sky soon. Opened enclosure, force filling of imager and spectro (174->134lb). Counterweights are 695 at balance. Run 6604, Supernova, 82-S 07:17 Slue to lambda=-159 7:21 goDrift lambda=-158. Focus is 170 at temperature around 5C. Skippy is happy from the beginning. In focus from frame 16. Skippy is good from the beginning. 7:32Z Start ramp frame 25, lambda=-156 At this time clouds go away and the weather conditions changed to almost photometric. Seeing is good and acceptable during whole stripe. 09:36Z End science, lambda=-125, frame 232. 9:46 Z End ramp and drift lambda=-122.5, frame 249. Segue stripe, 1406 N. 9:49Z slue to mu=99 Run 6605 9:59Z start rump mu=101 10:06 start science mu=103 In focus from the beginning (focus=-150) Skippy is happy from the beginning Seeing is good at the beginning. But seeing is bad from frame 139 mu 119.2. We did not noted any changes in the weather parameters. We continue 10 minuts more, seeing was still bad and we ended drift at 11:32Z frame 177, mu=125. It is a little time before twilight, only 44 minuts, Cart 4, plate 2386, seque 11:53 go-to-field, found stars 12:05 science exposure 1200. We deside to do calibration after science exposure as we want to use time before twi for science. We used first 11 min of exposure before twi, and 9 minuts after. Plate is very strange, the stars shifted to different sides of fibers and did not center on them even after our attemption to disable some fibers. We think about large rotator error. 12:28Z PostCalibration. We run sos_redo expnum=41530 to force the reduction. We obtained combined exposure on b1 32.1 and declared the plate as DONE. We had some minor problem this night: reboted photo-1 crate; turn off and on display photo 5 - stopped to display image imager cart had very low pressure when we open imager house, 26, server did not start mt window watcher complained about bad connection with server watcher complained about broken sftp connection with im camera, we reboot server and server did not start mt server window Cartridge 1 is mounted. 12:55 Telescope is closed. We wait twi to make twi PT flat, but it it is clouds again and we closed. 13:50Z EndNight is completed Spooling did not requested but Steph recommended us to do it all time. Spooling does not work, produce error message. [sn_user@sn10 ~]$ spool_photo_wrapper.cmd 6604 54064 ERROR: last run 6604 has no backup. Could not find /data/sn10.a/data/sndp_logs/6604.tar.gz . Could not find /data/sn10.a/data/sndp_logs/transfered/6604.tar.gz . Recommend backup before continuing. ***** ABORT ***** ---=== IMAGING RUN SUMMARY ===--- Run Time Stripe Lambda Last Flavor Comments Start End Begin End Frame ------------------------------------------------------------------------- 6601 21:56Z 22:09Z 100 O 102.05 105.08 34 ignore 6602 00:14Z 00:32Z 100 O 136.48 140.93 44 bias 6603 05:02Z 05:28Z 100 O -151.23 -144.83 38 bias - bad 6604 07:21Z 09:47Z 82 S -158.65 -122.13 258 science 6605 09:51Z 11:34Z 1406 N -85.61 -59.93 185 science ---=== IMAGING RUN DETAILS ===--- ---=== SKIPPY RESULTS ===--- Run Frame nFrames stars muErr muRms nuErr nuRms rot az el --------------------------------------------------------------------------- 6604 37 4 45 6.600 0.131 2.400 0.123 -0.01059 239 39 6604 90 4 45 6.200 0.148 2.200 0.111 -0.01015 239 39 6604 142 4 55 6.200 0.162 2.000 0.124 -0.00833 239 39 6604 194 4 52 5.900 0.140 1.800 0.110 -0.00861 239 39 6604 246 4 59 6.000 0.156 1.900 0.138 -0.00844 239 39 6605 12 4 0 0.000 0.000 0.000 0.000 -0.00000 0 0 6605 166 4 98 4.500 0.138 5.600 0.151 0.01842 216 73 ---=== LTMATCH RESULTS ===--- Run Field nFields alt az nGood rowMean rowSig colMean colSig rot ------------------------------------------------------------------------ 6604 48 1 38 301 15 -1.941 0.218 0.087 0.147 -0.00029 6604 48 1 38 301 20 -1.373 0.172 0.148 0.128 -0.00050 6604 48 1 38 301 13 -0.847 0.251 0.224 0.134 -0.00075 6604 48 1 38 301 24 -0.597 0.185 0.294 0.139 -0.00099 6604 48 1 38 301 9 -0.118 0.211 0.538 0.097 -0.00181 6604 66 1 38 301 13 -1.395 0.146 -0.505 0.312 0.00170 6604 66 1 38 301 18 -1.129 0.175 -0.236 0.254 0.00080 6604 66 1 38 301 11 -0.892 0.142 0.122 0.177 -0.00041 6604 66 1 38 301 11 -0.635 0.205 0.297 0.156 -0.00100 6604 66 1 38 301 17 -0.219 0.106 0.225 0.215 -0.00076 6604 101 1 38 301 17 -1.237 0.077 -0.439 0.204 0.00148 6604 101 1 38 301 12 -0.899 0.162 -0.324 0.200 0.00109 6604 101 1 38 301 16 -0.591 0.115 -0.240 0.160 0.00081 6604 101 1 38 301 14 -0.320 0.106 -0.119 0.141 0.00040 6604 101 1 38 301 17 -0.068 0.105 0.144 0.158 -0.00048 6604 119 1 38 301 18 -1.620 0.135 -0.725 0.116 0.00244 6604 119 1 38 301 15 -1.120 0.115 -0.562 0.120 0.00189 6604 119 1 38 301 15 -0.738 0.151 -0.287 0.247 0.00096 6604 119 1 38 301 15 -0.201 0.229 -0.137 0.116 0.00046 6604 119 1 38 301 6 0.176 0.224 0.193 0.127 -0.00065 6604 153 1 38 301 16 -1.455 0.137 -0.641 0.170 0.00215 6604 153 1 38 301 16 -0.951 0.163 -0.392 0.209 0.00132 6604 153 1 38 301 15 -0.410 0.353 -0.238 0.174 0.00080 6604 153 1 38 301 8 0.214 0.324 -0.103 0.174 0.00035 6604 153 1 38 301 20 0.598 0.223 0.049 0.193 -0.00017 6604 169 1 38 301 19 -2.033 0.259 -0.609 0.198 0.00205 6604 169 1 38 301 14 -1.683 0.255 -0.459 0.234 0.00154 6604 169 1 38 301 14 -1.468 0.148 -0.360 0.204 0.00121 6604 169 1 38 301 20 -0.875 0.216 -0.254 0.196 0.00085 6604 169 1 38 301 19 -0.264 0.144 -0.102 0.213 0.00034 6604 205 1 38 301 18 -1.324 0.267 -0.295 0.267 0.00099 6604 205 1 38 301 22 -0.894 0.216 -0.127 0.215 0.00043 6604 205 1 38 301 14 -0.442 0.231 0.154 0.270 -0.00052 6604 205 1 38 301 16 -0.155 0.266 0.368 0.183 -0.00124 6604 205 1 38 301 13 0.406 0.201 0.231 0.248 -0.00077 6604 221 1 38 301 18 -1.460 0.364 -0.802 0.201 0.00270 6604 221 1 38 301 15 -1.065 0.204 -0.584 0.320 0.00197 6604 221 1 38 301 20 -0.429 0.213 -0.457 0.325 0.00153 6604 221 1 38 301 16 0.092 0.147 -0.022 0.210 0.00007 6604 221 1 38 301 8 0.637 0.229 0.020 0.161 -0.00007 ---=== SPECTROSCOPY DATA SUMMARY ===--- Summary Checked (y/n): QA Procedures Done (y/n): UT Exp Time flavor comment (S/N)^2 totals ========================================== b1 r1 b2 r2 ----- sequence 41521, plate -9999 ------- 21:58 41521 0.0 bias ----- sequence 41522, plate 2386 ------- 22:08 41522 3.0 arc 22:10 41523 3.0 arc 22:14 41524 10.0 flat 22:16 41525 2.0 arc 23:00 41526 2.0 arc 23:02 41527 3.0 arc ----- sequence 41528, plate -9999 ------- 23:18 41528 0.0 bias 23:42 41529 0.0 bias ----- sequence 41530, plate 2386 ------- 12:27 41530 1201.0 target 12:30 41531 10.0 flat 12:33 41532 2.0 arc ---=== TELESCOPE OFFSETS AND SCALE I ===--- Time Instrument Az Alt Rot Scale pos offset pos offset pos offset ------------------------------------------------------------------------------ 07:39Z imager -59.13 0.0002 38.48 0.0034 43.66 -0.0097 1.000000 09:28Z imager -59.11 0.0002 38.48 0.0034 43.68 -0.0097 1.000000 10:43Z imager -28.72 0.0002 69.34 0.0034 39.93 -0.0097 1.000000 12:20Z 4 2386 50.36 0.0024 79.68 0.0028 226.01 0.0418 1.000170 ---=== TELESCOPE OFFSETS AND SCALE II ===--- ---=== FOCUS LOG ===--- setmir piston Temp Wind Time Inst scale M1 M2 Foc Az Alt (C) MPH Dir filt fwhm ------------------------------------------------------------------------------ 07:39Z imager 1.00000 0 -217 -174 -59 38.5 5.3 12 206 i23 1.09 09:28Z imager 1.00000 0 -199 -156 -59 38.5 4.1 12 214 i23 1.31 10:14Z imager 1.00000 0 -270 -142 -24 66.5 4.0 12 242 i23 10:43Z imager 1.00000 0 -264 -132 -29 69.3 4.0 10 266 i23 1.17 12:21Z 4 2386 1.00017 -1679 -221 50 50 79.7 3.2 8 254 -- 1.50 ---=== WEATHER LOG ===--- Wind Time Temp F Dewp F MPH Direction Dust DIMM Sky 20:48Z 53 21 14 227 (SW) 895 - clouds 21:20Z 51 21 13 200 (SSW) 920 - clouds 21:53Z 50 18 17 184 (S) 818 - 22:24Z 46 18 16 193 (SSW) 807 - 22:54Z 45 19 14 178 (S) 892 - 23:24Z 44 20 10 158 (SSE) 976 - 23:56Z 43 20 12 163 (SSE) 898 - 01:22Z 43 21 14 180 (S) 972 - 01:54Z 43 21 14 163 (SSE) 1014 - 02:26Z 43 21 15 178 (S) 1012 - 02:58Z 43 21 8 181 (S) 1032 - 03:31Z 43 20 17 183 (S) 1035 - 04:03Z 43 20 17 193 (SSW) 1110 - 04:35Z 43 20 15 187 (S) 997 - 05:08Z 43 19 14 188 (S) 973 - 05:38Z 43 18 16 205 (SSW) 1071 - 06:10Z 43 17 15 202 (SSW) 916 - 06:42Z 43 18 16 198 (SSW) 888 - clouds until this time 07:14Z 42 18 13 209 (SSW) 932 - 07:46Z 41 18 12 198 (SSW) 1090 - 08:20Z 41 18 14 212 (SSW) 1106 - 08:50Z 40 19 13 220 (SW) 1268 - 09:23Z 39 19 11 215 (SW) 1256 - 09:55Z 39 19 12 215 (SW) 1641 - 10:29Z 38 19 11 254 (WSW) 1866 - 10:59Z 38 18 11 260 (W) 721 - 11:30Z 38 18 10 261 (W) 752 - 12:01Z 38 18 10 259 (W) 905 - 12:32Z 37 19 8 257 (WSW) 1016 - 13:04Z 37 18 8 267 (W) 705 - 13:36Z 37 16 7 276 (W) 546 - 14:08Z 37 17 10 267 (W) 489 - ---=== TELESCOPE STATUS ===--- 23:45Z Doors opened, fans and blowers on, louvers opened. 06:45Z Enclosure off. 12:45Z Enclosure on, doors closed, fans & blowers off, louvers closed. 12:25 Telescope stowed at 30 degrees with cart 1 mounted. Counterweights at 250. Spectro dewar 13:30Z 111 lb, 26 psi; Imager dewar 13:30Z 120 lb, 20 psi ---=== SOFTWARE USED ===--- IOP/SOP: v4_28_1 Watcher: v2_35_0 MCP: v6_3_2 TPM: tpm_v3_5_0 AstroDa: v15_13 TCC: 2.7.2.1 sdssProcedures: v2_20 SoS: v5_1_4 hoggPT: v1_6_9 plate-mapper: v4_3_1 ---=== MIRROR NUMBERS ===--- PRIMARY: -------- Scale: 1.000000 MIGS TONIGHT NOMINAL Axial A 24.6760 24.6750 Axial B 26.2640 26.2540 Axial C 25.3290 25.3490 Trans D 13.5640 13.5760 Lateral E 9.9695 9.9568 Lateral F 11.5443 11.5316 GALILS Commanded: 4600. -2000. 2650. -9300. -9800. -9800. Actual: 4600. -2000. 2650. -9300. -9800. -9800. SETMIR VALUES PriDesOrient: 0.00 0.00 0.00 0.00 0.00 PriOrient: 0.00 0.00 0.00 0.00 0.00 SECONDARY: ---------- Focus: 0.00 Air Temp.: 7.2 Alt.: 30.032158 MIGS TONIGHT NOMINAL Axial A 1.5520 1.5550 Axial B 1.1350 1.1290 Axial C 1.1900 1.1910 Trans D -0.0530 -0.0530 GALILS Commanded: 1614005. 1594534. 1543076. -10550. -5800. Actual: 1614053. 1594719. 1543082. -10535. -5822. SETMIR VALUES SecDesOrient: 1257.00 20.00 -10.00 -100.00 176.64 SecOrient: 1256.94 19.95 -10.03 -100.46 176.59 ---=== PROBLEMS IN DETAIL ===--- tpm_TM_PMSSAvg 8.09786 tpm_TM_TelairAvg.VALA 9.315 sp1: Temp_Median 11 Temp_Hartmann_Bot 10 Temp_Hartmann_Top 10 Temp_Red_Cam_Bot 11 Temp_Red_Cam_Top 11 Temp_Blue_Cam_Bot 11 Temp_Blue_Cam_Top 11 sp2: Temp_Median 9 Temp_Hartmann_Bot 9 Temp_Hartmann_Top 9 Temp_Red_Cam_Bot 9 Temp_Red_Cam_Top 9 Temp_Blue_Cam_Bot 9 Temp_Blue_Cam_Top 9